OpenStudy (anonymous):

A net torque of magnitude 200 Nm is exerted on the rotor of an electric generator for 31 s. What is the magnitude of the angular mo- mentum of the rotor at the end of 31 s if it was initially at rest? Answer in units of kgm2/s

OpenStudy (anonymous):

are you given the radius ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

torque=dL/dt. torque*time=change in angular momentum (if torque is constant); thus the answer is 200*31 N*m*s. 1 N=1 kg*m/s^2; 1 N*m*s=1 kg*m^2/s. =200*31 kg*m^2/s

OpenStudy (anonymous):

Thank you :)